OpenStudy (anonymous):

3(x − 1) − (x + 5) = 2

OpenStudy (anonymous):

Hint: Distribute the 3. Then solve for x.

OpenStudy (ksaimouli):

dont forget the - sign

OpenStudy (ksaimouli):

be careful

OpenStudy (anonymous):

3(x − 1) − (x + 5) = 2 3x - 3 - x + 5 = 2 Now combine like terms.

OpenStudy (anonymous):

Can you finish it up?

OpenStudy (anonymous):

5

OpenStudy (anonymous):

3x - 3 - x + 5 = 2 2x + 2 = 2 2x = 0 x = 0

OpenStudy (anonymous):

@jazy i think you messed up distributing the second parenthesis. You needed to distribute the - sign, as if it were a - 1 in front of the parenthesis.

OpenStudy (anonymous):

3(x − 1) − (x + 5) = 2 3x -3 -x - 5 = 2 2x + 2 = 2 2x = 0 x = 0 oh x still equals 0.

OpenStudy (anonymous):

I forgot about that despite what @ksaimouli had said.

OpenStudy (anonymous):

Wait... 3(x − 1) − (x + 5) = 2 3x -3 -x - 5 = 2 ****-3 - 5 = -7**** 2x - 7 = 2 2x = 9 x = 4.5 So x = 4.5...@ilikequadraticsokay check to make sure I didn't mess up this time (:

OpenStudy (anonymous):

@ilikequadraticsokay

OpenStudy (anonymous):

3(x − 1) − (x + 5) = 2 3x -3 -x - 5 = 2 2x - 8 = 2 2x = 10 x = 5 That's what I got, I think you messed up with adding the - 5 and - 3